OpenStudy (anonymous):

what is x^5/x^3 simplified

OpenStudy (anonymous):

x^2 i guess xd

jimthompson5910 (jim_thompson5910):

When you divide variable expressions like this, you subtract the exponents of the terms with the same variable So \[\Large \frac{x^5}{x^3} = x^{5-3} = x^2\]